December 11, 2017

Service Status

Current Status 

Core network – Melbourne DC1 : OK

Wireless nodes – Major : OK

Wireless nodes – Micro : OK

Wireless nodes – Regional : OK

VoIP (voice) platform : OK

Upstream routing – Domestic : OK

Upstream routing – International : OK

Netflix : OK

YouTube : OK

Spotify : OK

Google : OK


Current known issues

Wednesday 24th November 2021 1525 AEDT
Customers connected to / through our Humevale (Whittlesea) node may experience degraded service.
Upstream provider issues have affected two optical fibre circuits into our network. Backup circuits are currently in place.
We will advise once we are provided additional information, we apologise for any inconvenience caused.
1708: Update: There is a confirmed Fibre Break at South Morang. Awaiting further details and OTDR testing on break location. Further updates to be provided in 60 minutes.
1757: Update: Carrier engineers have further isolated alarms to an Optus site in Thomastown VIC and our field engineer is on route and expected to arrive within the next 30 minutes.
1906: Update: Field engineer has arrived at Thomastown VIC and has begun testing the fibre to identify the distance to the break.
2005: Update: Fibre testing has completed and identified the break close to the South Morang POI. A fibre crew has been mobilised and currently en route to locate the fibre and assess damage, they will arrive in the next 30 – 40 minutes. The next update will be provided in 60 minutes.
2137: Update: The fibre tech has arrived at the POI and no obvious break is seen, and another test is being run on the fibre to isolate further.

We will provide the next update in 60 minutes.
2202: Update: The fibre technician has isolated the break upstream based on the latest results from testing the fibre and is heading to next closest fibre pit to continue efforts to isolate the break.

We will give our next update in 60 minutes.
2322: Update: The fibre engineer has located the break and found the cable has been damaged by construction work in the area. The damage to the fibre and conduit is extensive and will require an excavator to assist with exposing fibre and installing a new conduit. A civil crew is being mobilised now to head to site and assist with repair efforts.
Thu 25th Nov 0657: Update: Replacement Cable run and service restored.

Planned outages and works

Nil.

Recently resolved issues

2021

9th June 2021 1325 EST : Portions of our network were segmented in response to a flood of broadcast traffic from an upstream network partner. This resulted in a large number of changes to internet traffic paths (routes). We await a post incident review.
Thursday 10th June 2021 : Due to severe weather yesterday and overnight, some node sites remain without mains power. Battery backup has been performing well and we continue to monitor remaining capacity. Crews were dispatched today with generators and change-over batteries.
There may be delays attending to routine enquiries and faults while we manage this situation, we thank you for your patience.
Tuesday 13th July 2021 11:40am – 9:51pm : Degraded service surrounding our Humevale node due to an upstream carrier snapped optical fibre in Thomastown. New cable hauled and spliced.
Fri 23rd July 2021 7:30pm AEST
A power controller located at our Broadford (Mt Hickey) site failed. Engineers replaced the unit.
Tuesday 14th September 2021 1015 AEST : A power surge at our Ivanhoe tower site resulted in a failed power controller. Engineers attended with replacement parts.
Thursday 28th October 2021 : Electrical storms and extreme winds caused various issues. Parts were replaced and backup power installed.